OpenStudy (anonymous):

Multiply and simplify: (4y − 1)(8y + 3) (A) 32y2 − 4y − 3 (B) 32y2 + 4y + 3 (C) 32y2 − 4y + 3 (D) 32y2 + 4y − 3

OpenStudy (anonymous):

4y(8y+3)-1(8y+3) =32y2+12y-8y-3 simplify it nw..

OpenStudy (anonymous):

u got the ans?

OpenStudy (anonymous):

Not really...

OpenStudy (anonymous):

add similar terms...

OpenStudy (anonymous):

for ex. 3y2-2y2+4y-2y =y2+2y in the way...try to add simillar terms...

OpenStudy (anonymous):

Oh okay, thanks (:

OpenStudy (anonymous):

wat ans u got?

OpenStudy (anonymous):